OpenStudy (anonymous):

The scores listed below are the results from students taking a dexterity test. What is the 75th percentile of these scores? 1, 1, 1, 1.5, 2, 2, 2, 2.5, 3 2 2.25 2.5

OpenStudy (anonymous):

@texaschic101 Will u help me

OpenStudy (anonymous):

Will u please tell me how u got that answer

OpenStudy (texaschic101):

of course.....do you know how to find quartiles ?

OpenStudy (anonymous):

Idk

OpenStudy (texaschic101):

Lets start again.....do you know how to find the median (middle number ) of this set of data ?

OpenStudy (anonymous):

Oh ya

OpenStudy (texaschic101):

what is it ?

OpenStudy (anonymous):

10

OpenStudy (texaschic101):

no....the middle number...there is no 10 in your data. How many numbers are there ?

OpenStudy (anonymous):

nvm its 2

OpenStudy (texaschic101):

yes...that is the middle number also called quartile 2 or the 50th percentile. Now look at the numbers before the median....1,1,1,1.5....take the median of this....since there is two medians we need to average....(1 + 1)/2 = 2/2 = 1. That 1 is the 25th percentile or quartile 1. You need to do the same for the numbers after the median....2,2,2.5,3....can you find the middle term in these numbers ? Because that will be quartile 3, also known as the 75th percentile. I was wrong about the answer....I was counting the numbers wrong...I didn't see the decimals. 2 is not the answer

OpenStudy (texaschic101):

do you understand ?

OpenStudy (texaschic101):

take the numbers after the median...2,2,2.5,3.....find the middle number. Since there is two middle numbers, we average them...(2 + 2.5)/2 = 4.5/2 = 2.25. 2.25 is your 75th percentile or quartile 3.

OpenStudy (anonymous):

So is it 2.25 or 2